OpenStudy (anonymous):

Which of the following is the slope between the two points (-5, -1) and (6, -4)?

OpenStudy (anonymous):

The slope between two points is equal to the difference in their y values minus the difference in their x values - \[\Delta y / \Delta x\] It doesn't matter which point you use first, but you need to be consistent. So I'll use the point (6, -4). The x value of this point is 6, the y value is -4. So the slope is then: Change in y = -4 - -1 = -4+1=-3 Change in x = 6 - -5 = 6+5=11 So the slope is -3/11.

OpenStudy (anonymous):

Consider the line that passes through the points (4, -3) and (2, -1). Below are two different equations. Identify the true statement concerning both equations. Equation #1 Equation #2 y - 3 = -1(x + 4) y - 1 = -1(x + 2)

OpenStudy (anonymous):

@vinnv226

OpenStudy (anonymous):

^ I'm sorry, in my first reply I said "minus" when I meant "divided by" in the first line ^

OpenStudy (anonymous):

ohh okay ")@vinnv226

OpenStudy (anonymous):

Could you explain what this new question is looking for? Is there anything more to the problem? I would guess you're being asked to find the equation of the line that passes through those two points but I'm a little confused by the wording you used.

OpenStudy (anonymous):

Thatss all it tells me to do :oo

OpenStudy (anonymous):

@vinnv226

OpenStudy (anonymous):

I'm not quite sure what the question is asking. If you distribute and solve each of those two equations you'll find that they, in fact, are the same graph.

OpenStudy (anonymous):

im i lovejusti25 i got banned lmao @vinnv226

OpenStudy (anonymous):

Which of the following is the slope between the two points (1, 5) and (6, 4)? @vinnv226

OpenStudy (anonymous):

This is solved the same way as the original problem, that is, by dividing the change in y by the change in x. Change in y = 4-5=-1 Change in x= 6-1=5 So the slope is -1/5.

OpenStudy (anonymous):

can you help with more ? I have 3 tests to take that depend on my grade -.- @vinnv226

OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (2, 4) and has a slope of 3.

OpenStudy (anonymous):

@vinnv226

OpenStudy (anonymous):

One way to represent a line is the following equation: \[y - y _{1} = m(x-x_{1})\] Where x1 and y1 are the coordinates of a point on that line, and m is the slope. So, since you know the slope is 3 and a point on the line is (2,4), you can plug in m=3, x1=2 and y1=4, and get: \[y-4 = 3(x-2)\] And now you just need to simplify.

OpenStudy (anonymous):

idk how @vinnv226

OpenStudy (anonymous):

Well, often when we simplify, we get it in a y= form. So, on the right side, we'll distribute the 3: \[y-4=3x-6\] And then we add 4 to both sides, and we're done: \[y=3x-2\]

OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (6, 4) and is parallel to y =1/3 x - 3. @vinnv226

OpenStudy (anonymous):

For two lines to be parallel, they must have the same slope. So the slope of our answer is going to be 1/3. Then, since we have the point (6,4) we can use the same equation we used above and plug in: \[y-4=(1/3)(x-6)\]Try practicing simplifying these types of equations to get them in y= forms.

OpenStudy (anonymous):

its hard @vinnv226